Please solve this inequality
2(3-2m)is greater than or equal to 8m+3
Thank you

Answer by checkley71(8403)

2(3-2M)>=8M+3
6-4M>=8M+3
-4M-8M>=3-6
-12M>=-3
M<=-3/-12 DIVIDING BY A NEGATIVE REVERSES THE INEQUALITY SIGN.
M<1/4
LET M=1/8 FOR THE PROOF
2(3-2*1/8>=8*1/8+3
2(3-1/4)>=1+3
2(2.75)>=4
5.5>=4